Describe verbally how to solve y mx c = + . What assumptions have you made about the value of m ? 


1
Expert's answer
2021-10-13T13:48:45-0400

One form of the equation of a straight line is called the slope-intercept form because it contains information about these two properties.


The equation of a straight line:


Any equation of the form "y = mx + c" where m and c are fixed numbers, (i.e. constants), has a graph which is a straight line.

For example, "y = 3x + 5, y = \frac{2}{3} x + 8" and "y = \u22123x \u2212 7" all have graphs which are straight lines.